OpenStudy (chycora):

2. (1 pt) Solve the word problem. The height of a triangle with base b and area A is given by the formula h = . What is the height of a triangle with area 26 cm2 and base 8 cm? A. 1.625 cm B. 3.25 cm C. 3.5 cm D. 6.5 cm

OpenStudy (cggurumanjunath):

26/8

OpenStudy (cggurumanjunath):

h = A / b = 26 / 8 = 3.25 cm

OpenStudy (anonymous):

Yea its B

OpenStudy (cggurumanjunath):

d

OpenStudy (chycora):

so it's d or b

OpenStudy (cggurumanjunath):

A = 0.5 * b * h 26 = 0.5 *8 * h

OpenStudy (cggurumanjunath):

26 = 4 * h h = 26 / 4 = 13 / 2 = 6.5 cm

OpenStudy (chycora):

so d thanx

OpenStudy (cggurumanjunath):

Sorry it is D

OpenStudy (chycora):

ur kool thanx